OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 6. x y 1 2 0 2 3 -10

OpenStudy (anonymous):

A. y=-2x^2+2x-2; -58 B. y=2x^2-2x-2; 60 C. y=2x2-2x-2; 58 D. y=-2x2+2; -58

OpenStudy (cwrw238):

those values don't fit any of the above eg x = 0, y =2 fits D but x = 1 y =2 does not x=0, y=2 does not fit A,B or C